Can someone help me with my project? Here is the problem.

Cyber Systems, Inc. reported the following shareholders’ equity: Shareholders’ equity: Preferred shares, $2.00 Authorized - 1,000,000 shares Issued 150,000 shares $ 3,750,000 Common shares Authorized - 5,000,000 shares Issued - 800,000 shares 6,800,000 Retained earnings 6,855,180 Total shareholders’ equity $17,405,180

Required: 1. What was the average issue price per share of the common shares?

2. What was the average issue price per share of the preferred shares?

3. Assume that net income for the year was $825,000 and that issued shares of both common and preferred shares remained constant during the year. Journalize the entry to close net income to Retained Earnings. What was the amount of earnings per share?

4. Journalize the issuance of 10,000 additional common shares at $22.50 per share. Use the same account titles as shown in the problem.

5. How many common shares are outstanding after the 10,000 additional shares have been sold?

6. How many common shares would be outstanding after the corporation split its common shares 2 for l? What is the new average issue price?

7. Journalize the declaration of a stock dividend when the market price of the common shares is $22.50 per share. Consider each of the following stock dividends independently:

a. Cyber Systems, Inc., declares a 10% common stock dividend on shares outstanding after the 2- for-l split.

b. Determine the average issue price per common share after the distribution of the stock dividend.

8. Journalize the following common shares transactions in the order given: a. Cyber Systems, Inc., repurchases 2,500 of its common shares at $25 per share. (HINT: Need new average issue price from 7.b) b. One month later, the corporation sells 1,000 of its common shares for $27 per share. c. An additional 1,000 of its common shares are sold for $22 per share.

9. Prepare an updated (after considering all the above) Statement of Shareholders’ Equity and Statement of Retained Earnings for Cyber Systems, Inc. (HINT: Use T-accounts to show your work/calculations for preferred shares, common shares, retained earnings and number of common shares).
Answer #1

1. Average issue price per share of common share = gross proceeds / no. of shares issued

= 6800000/800000 = 8.5 per share

2. Average issue price per share of preferred share = gross proceeds / no. of shares issued

= 3750000/ 150000 = 25 per share

3. journal entry

sr. no. particulars debit credit
1. net income $ 825000
retained earnings $ 825000

4.journal entry

particulars debit credit
bank 225000
common shares 225000
(22.5*10000)

since no paid in capital is created in the balance sheet, whole receipt from issue of shares is recorded in common share isued account.

5.

no. of shares
issued shares at the beginning of the year 800000
add: new shares issued 10000
total shares outstanding 810000

6. comon shares outstanding after split off of the shares 2 for 1 = 810000*2 = 1620000

new average share price = (6800000+225000)/1620000 = 4.34 (round off)